# Learning Multimodal Embeddings for Traffic Accident Prediction and Causal Estimation

### v2

@@ B.2. Omitted Experiment Results
In addition to AUROC, we also evaluate the performance of different baselines using the precision and recall scores as additional metrics.
@@
Causal analysis results of PSM and DR.
To enhance the causal analysis, we add Propensity Score Matching (PSM) and Doubly Robust (DR) estimators for more reliable treatment effect estimation, using one nearest neighbor and evaluating effects by road type. The results show a clear increase in accident risk for motorway segments, with effects of around $20\%$. To assess robustness, we also vary the number of nearest neighbors in the matching step. The estimates stay consistent across settings, showing that the DR estimator is stable.
@@ Appendix C Privacy Details and Leakage Audit
All data sources used in this study are publicly available and released under open data licenses by the corresponding agencies. State Departments of Transportation provide traffic accident and traffic volume records through official open data portals that support academic and non-commercial use. The OSMnx Street Network Dataverse is distributed under the Open Database License (ODbL), which allows reuse under standard attribution and share-alike terms. Mapbox imagery is obtained through the official Static Tiles API under the standard developer terms of service. Weather data is collected from the official Meteostat API, which offers open access to historical and near-real-time meteorological observations for research. Table 10 lists all sources, license statements, and links.
All datasets used in this work contain only anonymized records released by public agencies. No record includes personal identifiers, and no field can be linked back to individuals. The spatial resolution of the satellite and grid-level imagery does not allow identification of people, vehicles, or private property details. Our processing pipeline uses only node-level and segment-level information derived from public road networks, which ensures that the final dataset remains fully anonymous and consistent with public data release standards.
In addition to these safeguards, we conduct a full leakage audit to verify that no hidden pathways for re-identification remain. We check that accident counts are aggregated at the segment or monthly level, that imagery does not contain fine-grained visual cues tied to individuals, and that no timestamps, device traces, or location histories enter the dataset during pre-processing. We also confirm that structural information from OSMnx cannot be cross-matched with any external source to recover personal movement patterns. These checks show that the final dataset cannot be used to infer identities or sensitive attributes. In our experiments, we divide the entire time series into three non-overlapping segments for training, validation, and testing, as shown in Table 8. The splits are strictly chronological, ensuring that no future information leaks into the training or validation sets.
Table 10. Summary of the data sources used in our dataset construction, including links to official repositories for traffic accident records, traffic volume statistics, and APIs to road network data, weather observations, and satellite imagery. They are official records and serve as the foundation for building our multimodal traffic safety dataset.
@@
Satellite Image
MapBox API https://console.mapbox.com/
Road Networks

### v3

@@ B.2. Omitted Experiment Results
In addition to AUROC, we evaluate the performance of different baselines using precision and recall as additional metrics.
@@
Causal analysis results of PSM and DR
To enhance the causal analysis, we add Propensity Score Matching (PSM) and Doubly Robust (DR) estimators for more reliable treatment effect estimation, using one nearest neighbor and evaluating effects by road type. The results show a clear increase in accident risk for motorway segments, with effects of around $20\%$. To demonstrate consistent treatment effects rather than sensitivity to a specific matching setup, we also vary the number of nearest neighbors in the matching step. The estimates stay consistent across settings, showing that the DR estimator is stable.
@@ Appendix C Privacy Details and Leakage Audit
All data sources used in this study are publicly available and released under open data licenses by the corresponding agencies. State Departments of Transportation provide traffic accident and traffic volume records through official open data portals that support academic and non-commercial use. The OSMnx Street Network Dataverse is licensed under the Open Database License (ODbL), which allows reuse under standard attribution and share-alike terms. Mapbox imagery is sourced through the official Static Tiles API under the standard developer terms of service. Weather data is collected from the official Meteostat API, which offers open access to historical and near-real-time meteorological observations for research. Table 10 lists all sources, license statements, and links.
All datasets used in this work contain only anonymized records released by public agencies. No record includes personal identifiers, and no field can be linked back to individuals. The spatial resolution of the satellite and grid-level imagery does not allow identification of people, vehicles, or private property details. Our processing pipeline uses only node- and segment-level information derived from public road networks, ensuring that the final dataset remains fully anonymous and consistent with public data release standards.
In addition to these safeguards, we conduct a full leakage audit to verify that no hidden pathways for re-identification remain. We check that accident counts are aggregated at the segment or monthly level, that imagery does not contain fine-grained visual cues tied to individuals, and that no timestamps, device traces, or location histories are included in the dataset during pre-processing. We also confirm that structural information from OSMnx cannot be cross-matched with any external source to recover personal movement patterns. These checks show that the final dataset cannot be used to infer identities or sensitive attributes. In our experiments, we divide the entire time series into three non-overlapping segments for training, validation, and testing, as shown in Table 8. The splits are strictly chronological, ensuring that no future information leaks into the training or validation sets.
Table 10. Summary of the data sources used in our dataset construction, including links to official repositories for traffic accident records, traffic volume statistics, and APIs to road network data, weather observations, and satellite imagery. They are official records and form the foundation of our multimodal traffic safety dataset.